# Good Seed Makes a Good Crop: Discovering Secret Seeds in Text-to-Image Diffusion Models

### v1

@@ 4.1 High-Fidelity Inference
In Section 3.4, we observed that ‘golden’ seeds tend to generate images with significantly better quality and human preference alignment. This inspires us to think—how much can we improve the image quality compared to random generations by simply leveraging these ‘golden’ seeds?
Specifically, we identified $k$ ‘golden’ seeds that excel in both image quality and human preference alignment. We subsequently tested these $k$ ‘golden’ seeds by generating images with a different set of 10,000 prompts to evaluate their performance relative to random seeds. We identified $k=65$ ‘golden’ seeds for SD 2.0 and $k=67$ for SDXL Turbo, where $k$ was determined by selecting seeds that ranked among the top 256 in both FID [40] and HPS v2 [52]. We propose that a sampling pool of 60+ ‘golden’ seeds is sufficiently large in practical applications for a single prompt. As demonstrated in Table 1, leveraging these well-chosen seeds significantly improves the FID and HPS v2 scores for both SD 2.0 and SDXL Turbo, and on both MS-COCO [24] and the PartiPrompts benchmark [56].
@@
4.2 Controllable Diversity in Style and Composition
A typical image generation interface presents the user with four samples per prompt. Moreover, prior methods aim to promote the diversity of generated images using primarily gradient-based methods, such as Particle Guidance [9]. In Section 3.4, our results highlight that the choice of seed has a strong influence on the stylistic and spatial attributes in the generated images. Therefore, we explore whether we can obtain more diverse images in style or composition by merely sampling ‘diverse’ seeds.
To select four diverse seeds per prompt, we represent each seed by a feature vector capturing its style or composition, as discussed in Section 3.4. We then employ farthest point sampling using these features. Specifically, we randomly pick the first seed $s_{0}\sim\mathcal{U}\{0,1023\}$ and iteratively select the next three seeds to maximize the distance in feature space from the already selected seeds.
@@ 4.2 Controllable Diversity in Style and Composition
where $S$ is our set of diverse seeds. To evaluate whether our well-chosen seeds improve diversity over random seeds and Particle Guidance [9], we calculate the similarity between images synthesized from a different set of $P$ prompts, where $P=500$ LLaVA [25] dense captions for image style and $P=440$ synthetic prompts for image composition. In particular, we measure the pairwise cosine similarity of image features and average the similarity scores across prompts. Intuitively, a lower pairwise similarity score means higher diversity. Mathematically, the metric score is represented as:
@@
where there are $P$ prompts and $\mathbf{f}$ denotes the feature vector representing image style or composition. We typically use $C=4$ images per prompt, but it’s important to note that if no objects are detected in an image, then the image is not used to compute similarity. In Table 2, we observe that our diverse seeds outperform random seeds and Particle Guidance [9] in generating images with varying styles and compositions for SD 2.0. Interestingly, our well-chosen seeds aid in diversifying image composition for SD 2.0 but not for SDXL Turbo. We show visual comparisons in Figure 11.
Figure 11: We show that simply generating images using "diverse" seeds can promote more variation in style (top) and image composition, as measured by object centroid, size, and depth (bottom).
@@
Style Similarity ()($\downarrow$)( ↓ ) Composition Similarity ()($\downarrow$)( ↓ ) Style Similarity ()($\downarrow$)( ↓ ) Composition Similarity ()($\downarrow$)( ↓ )
Random Seeds 0.981 $\pm$ 0.001 0.971 $\pm$ 0.001 0.993 $\pm$ 0.000 0.988 $\pm$ 0.000
Particle Guidance 0.980 $\pm$ 0.000 0.972 $\pm$ 0.000
Our Diverse Seeds 0.970 $\pm$ 0.000 0.961 $\pm$ 0.001 0.984 $\pm$ 0.000 0.988 $\pm$ 0.000
@@ 4.3 Improved Text-based Inpainting
In Sections 4.1 and 4.2, we demonstrated that carefully selecting the seed provides a straightforward, training-free approach to enhance the visual quality, human preference, and diversity of images generated by text-to-image diffusion models. But, the potential of image generation extends beyond text-to-image applications. This poses an intriguing question—can we also uncover ‘golden’ seeds for text-based image inpainting tasks, such as object removal and object completion?
As described in Section 3.2, we gathered 500 pairs of images and inpainting masks for the object removal and object completion applications. We employed the text prompt "clear background" for the removal case, and we used a prompt corresponding to the original object category for the completion case. We then generated images using a text-based diffusion inpainting model. We observed that some images contain unwanted text in the inpainting region that often mimics the prompt. To quantify the presence of text, we applied optical character recognition [16] and calculated the average proportion of text artifacts within the inpainting mask across all images from each seed. As illustrated in Figure 12, certain seeds are prone to inserting text in both removal and completion scenarios.
Figure 12: We discover that certain seeds tend to insert unwanted text within the inpainting region, outlined in pink. Top: We aim to remove the object using the prompt "clear background." Bottom: We attempt to complete the object using a prompt that specifies the object category.
5 Conclusion, Limitations, and Broader Impacts
In this work, we investigated the role of "random" seeds in the reverse diffusion process, exploring their differentiability based on generated images and their impact on interpretable visual dimensions. Notably, our 1,024-way classifier trained to predict the seed number for a generated image achieved over 99.9% test accuracy in just a few epochs. Encouraged by this finding, we conducted extensive analyses and identified ‘golden’ seeds that consistently produce images with better visual quality and human preference alignment. Additionally, we discovered that certain seeds create ‘grayscale’ images, add borders, or insert text during inpainting. Our studies also show that seeds influence the image composition, affecting object position, size, and depth. Leveraging these insights, we propose downstream applications such as high-fidelity inference and diversified generation for text-to-image diffusion models by merely sampling these special seeds. Our analyses offer new perspectives on enhancing image synthesis during inference without significant computational overhead.
Moreover, it’s important to recognize that we employ text-to-image diffusion models pretrained on large-scale, uncurated web data that may contain biases and errors, and our text prompts include dense captions of MS-COCO [24] images that may produce human imagery. Additionally, due to budget constraints, we primarily study the impact of 1,024 seeds out of all possible seed values.
@@ Appendix A Data Generation
In Section 3.2, we employed pretrained model checkpoints and implementations from the Hugging Face diffusers library [50]. Specifically, for text-to-image generation, we used Stable Diffusion 2.0 ("stabilityai/stable-diffusion-2-base") with a DDIM scheduler, and SDXL Turbo ("stabilityai/sdxl-turbo"). For text-based image inpainting, we utilized the SD 2.0 inpainting model ("stabilityai/stable-diffusion-2-inpainting"). Furthermore, our 1,024 seeds range from 0 to 1,023 inclusive, and we use torch.Generator("cuda").manual_seed(seed) to assign the seed used by the model.
@@ A.1 Synthetic Prompts for Image Composition Analysis
We create a set of 880 prompts by pairing 40 object categories with 22 modifiers in the format "a [modifier] [object category]". These modifiers include 21 adjectives and the empty string.
@@ A.2 Dataset for Inpainting Applications
We curated 500 pairs of images and inpainting masks for object removal and object completion applications, as described in Section 3.2. In particular, for the object removal use case, we employed images and annotations from the Open Images dataset [19, 21], and we used "clear background" as the text prompt. To create the inpainting mask, we dilated the instance segmentation mask to ensure coverage of the object. Additionally, for the object completion use case, we sampled images from the MS-COCO dataset [24] and used InstaOrder [22] to determine occlusion relationships to create inpainting masks. We used the category of the object to complete as the text prompt.
@@ A.3 Licenses for Existing Datasets
The MS-COCO dataset [24] and the PartiPrompts benchmark [56] are under a CC BY 4.0 license. Additionally, for the Open Images dataset [19, 21], the images are under a CC BY 2.0 license and the annotations are under a CC BY 4.0 license.
@@ Appendix B Classifier for Predicting Seed Number
We trained a lightweight transformer, EfficientFormer-L3 [23], to predict the seed used to generate an image. For our 1,024-way classification task, we utilized 9,000 training, 1,000 validation, and 1,000 test images per seed as mentioned in Section 3.3. The prompts for these images are dense captions by LLaVA 1.5 [25]. Moreover, we set a batch size of 128 and train for six epochs, which obtains a model checkpoint with over 99.9% validation and test accuracy. Our classifier uses the AdamW optimizer [26] with learning rate 0.0002 and weight decay 0.05. We apply data augmentations during training, which include resizing each image to have a shorter edge of size $224$ using bicubic interpolation, center cropping the image to size $224\times 224$, and randomly flipping the image horizontally with probability $0.5$. During validation and testing, we only resize and center crop the images.
@@ Appendix C Compute Resources
To generate our dataset in Section 3.2, we utilized 32 A100 GPUs for roughly 24 days. Additionally, all the experiments in Sections 3.3, 3.4, and 4 were performed on an RTX 4090 GPU with 24GB of memory. One of the longest experiments was training the classifier to predict seed number in Section 3.3, which took at most three days.
@@ Appendix D Additional Qualitative Results
We provide extra visualizations of the Grad-CAM from our classifier that predicts seed number in Figure 13. We also show more examples of seeds that often produce a ‘border’ around the image in Figure 14. Moreover, we present additional examples of good seeds and seeds that generate "text artifacts" for object removal and completion applications in Figures 15 and 16, respectively.

### v2

@@ 4.1 High-Fidelity Inference
In Sec. 3.4, we observed that ‘golden’ seeds tend to generate images with significantly better quality and human preference alignment. This inspires us to think—how much can we improve the image quality compared to random generations by simply leveraging these ‘golden’ seeds?
Specifically, we identified $k$ ‘golden’ seeds that excel in both image quality and human preference alignment. We subsequently tested these $k$ ‘golden’ seeds by generating images with a different set of 10,000 prompts to evaluate their performance relative to random seeds. We identified $k=65$ ‘golden’ seeds for SD 2.0 and $k=67$ for SDXL Turbo, where $k$ was determined by selecting seeds that ranked among the top 256 in both FID [41] and HPS v2 [53]. We propose that a sampling pool of 60+ ‘golden’ seeds is sufficiently large in practical applications for a single prompt. As demonstrated in Tab. 1, leveraging these well-chosen seeds significantly improves the FID and HPS v2 scores for both SD 2.0 and SDXL Turbo, and on both MS-COCO [25] and the PartiPrompts benchmark [57].
@@
4.2 Controlling Diversity in Style and Composition
A typical image generation interface presents the user with four samples per prompt. Moreover, prior methods aim to promote the diversity of generated images using primarily gradient-based methods, such as Particle Guidance [9]. In Sec. 3.4, our results highlight that the choice of seed has a strong influence on the stylistic and spatial attributes in the generated images. Therefore, we explore whether we can obtain more diverse images in style or composition by merely sampling ‘diverse’ seeds.
To select four diverse seeds per prompt, we represent each seed by a feature vector capturing its style or composition, as discussed in Sec. 3.4. We then employ farthest point sampling using these features. Specifically, we randomly pick the first seed $s_{0}\sim\mathcal{U}\{0,1023\}$ and iteratively select the next three seeds to maximize the distance in feature space from the already selected seeds.
@@ 4.2 Controlling Diversity in Style and Composition
where $S$ is our set of diverse seeds. To evaluate whether our well-chosen seeds improve diversity over random seeds and Particle Guidance [9], we calculate the similarity between images synthesized from a different set of $P$ prompts, where $P=500$ LLaVA [26] dense captions for image style and $P=440$ synthetic prompts for image composition. Specifically, we measure the pairwise cosine similarity of image features and average the similarity scores across prompts. A lower pairwise similarity score means higher diversity. Mathematically, the similarity is computed as:
@@
where there are $P$ prompts and $\mathbf{f}$ denotes the feature vector representing image style or composition. We typically use $C=4$ images per prompt, but it’s important to note that if no objects are detected in an image, then the image is not used to compute similarity. In Tab. 2, we observe that our diverse seeds outperform random seeds and Particle Guidance [9] in generating images with varying styles and compositions for SD 2.0. Interestingly, our well-chosen seeds aid in diversifying image composition for SD 2.0 but not for SDXL Turbo. We show comparisons in Fig. 11.
Figure 11: We show that simply generating images using “diverse" seeds can promote more variation in style (top) and image composition, as measured by object centroid, size, and depth (bottom).
@@
Style ()($\downarrow$)( ↓ ) Composition ()($\downarrow$)( ↓ ) Style ()($\downarrow$)( ↓ ) Composition ()($\downarrow$)( ↓ )
Random Seeds 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00
Particle Guidance 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00
Our Diverse Seeds 0.97 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.99 $\pm$ 0.00
Figure 12: Certain seeds tend to insert unwanted text within the inpainting region, outlined in pink. Top: We aim to remove the object using the prompt “clear background." Bottom: We attempt to complete the object using a prompt that specifies the object category.
@@ 4.3 Improved Text-based Inpainting
In Sec. 4.1 and 4.2, we showed that carefully selecting the seed provides a straightforward, training-free approach to enhance the visual quality, human preference, and diversity of images generated by text-to-image diffusion models. But, the potential of image generation extends beyond text-to-image applications. This poses an intriguing question—can we also uncover ‘golden’ seeds for text-based image inpainting tasks, such as object removal and completion?
As described in Sec. 3.2, we gathered 500 pairs of images and inpainting masks for the removal and completion applications based on the Open Images dataset [22, 20]. We employed the prompt “clear background" for the removal case, and we used a prompt corresponding to the original object category for the completion case. We then generated images using a text-based diffusion inpainting model. We observed that some images contain unwanted text in the inpainting region that often mimics the prompt. To quantify the presence of text, we applied optical character recognition [17] and computed the mean proportion of text artifacts within the inpainting mask across all images from each seed. As shown in Fig. 12, certain seeds tend to insert text in both removal and completion scenarios.
5 Conclusion
In this work, we investigated the role of “random" seeds in the reverse diffusion process, exploring their differentiability based on generated images and their impact on interpretable visual dimensions. Notably, our 1,024-way classifier trained to predict the seed number for a generated image achieved over 99.9% test accuracy in just a few epochs. Encouraged by this finding, we conducted extensive analyses and identified ‘golden’ seeds that consistently produce images with better visual quality and human preference alignment. Additionally, we discovered that certain seeds create ‘grayscale’ images, add borders, or insert text during inpainting. Our studies also show that seeds influence the image composition, affecting object position, size, and depth. Leveraging these insights, we propose downstream applications such as high-fidelity inference and diversified generation for text-to-image diffusion models by merely sampling these special seeds. Our analyses offer new perspectives on enhancing image synthesis during inference without significant computational overhead.
Acknowledgements. This work is supported by funds provided by the National Science Foundation and by DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence).
@@ Appendix A Data Generation
In Sec. 3.2, we employed pretrained model checkpoints and implementations from the Hugging Face diffusers library [51]. Specifically, for text-to-image generation, we used Stable Diffusion 2.0 (“stabilityai/stable-diffusion-2-base") with a DDIM scheduler, and SDXL Turbo (“stabilityai/sdxl-turbo"). For text-based image inpainting, we utilized the SD 2.0 inpainting model (“stabilityai/stable-diffusion-2-inpainting"). Furthermore, our 1,024 seeds range from 0 to 1,023 inclusive, and we use torch.Generator("cuda").manual_seed(seed) to assign the seed used by the model.
@@ A.1 Synthetic Prompts for Image Composition Analysis
We create a set of 880 prompts by pairing 40 object categories with 22 modifiers in the format “a [modifier] [object category]". These modifiers include 21 adjectives and the empty string.
@@ A.2 Dataset for Inpainting Applications
We curated 500 pairs of images and inpainting masks for object removal and object completion applications, as described in Sec. 3.2. In particular, for the object removal use case, we employed images and annotations from the Open Images dataset [20, 22], and we used “clear background" as the text prompt. To create the inpainting mask, we dilated the instance segmentation mask to ensure coverage of the object. Additionally, for the object completion use case, we sampled images from the MS-COCO dataset [25] and used InstaOrder [23] to determine occlusion relationships to create inpainting masks. We used the category of the object to complete as the text prompt.
@@ A.3 Licenses for Existing Datasets
The MS-COCO dataset [25] and the PartiPrompts benchmark [57] are under a CC BY 4.0 license. For the Open Images dataset [20, 22], the images are under a CC BY 2.0 license and the annotations are under a CC BY 4.0 license.
@@ Appendix B Classifier for Predicting Seed Number
We trained a lightweight transformer, EfficientFormer-L3 [24], to predict the seed used to generate an image. For our 1,024-way classification task, we utilized 9,000 training, 1,000 validation, and 1,000 test images per seed as mentioned in Sec. 3.3. The prompts for these images are dense captions by LLaVA 1.5 [26]. Moreover, we set a batch size of 128 and train for six epochs, which obtains a model checkpoint with over 99.9% validation and test accuracy. Our classifier uses the AdamW optimizer [27] with learning rate 0.0002 and weight decay 0.05. We apply data augmentations during training, which include resizing each image to have a shorter edge of size $224$ using bicubic interpolation, center cropping the image to size $224\times 224$, and randomly flipping the image horizontally with probability $0.5$. During validation and testing, we only resize and center crop the images.
@@ Appendix C Compute Resources
To generate our dataset in Sec. 3.2, we utilized 32 A100 GPUs for roughly 24 days. Additionally, all the experiments in Sec. 3.3, 3.4, and 4 were performed on an RTX 4090 GPU with 24GB of memory. One of the longest experiments was training the classifier to predict seed number in Sec. 3.3, which took at most three days.
@@ Appendix D Additional Qualitative Results
We provide extra visualizations of the Grad-CAM from our classifier that predicts seed number in Figure 13 of the supplemental. We also show more examples of seeds that often produce a ‘border’ around the image in Figure 14 of the supplemental. Moreover, we present additional examples of good seeds and seeds that generate “text artifacts" for object removal and completion applications in Figures 15 and 16 of the supplemental, respectively.